# Tunneling in double-well potentials within stochastic quantization: Application to ammonia inversion

### v1

@@ Abstract
Stochastic quantization—introduced by Nelson in 1966—describes quantum behavior as a conservative diffusion process in which a particle undergoes Brownian-like motion with a fluctuation amplitude set by Planck’s constant. While it fully reproduces conventional quantum mechanics, this approach provides an alternative framework that enables the study of dynamical quantities not easily defined within the standard formulation. In the present work, stochastic quantization is employed to investigate tunneling-time statistics for bound states in double-well potentials. Using first-passage time theory within the stochastic quantization framework, both the mean tunneling time, $\bar{\tau}$, and the full probability distribution, $p({\tau})$, are computed, and the theoretical predictions are validated through extensive numerical simulations of stochastic trajectories for the two potentials considered as representative cases. For the square double-well potential, analytical expressions for $\bar{\tau}$ are derived and show excellent agreement with simulations. In the high-barrier limit, the results reveal a direct relation between the stochastic-mechanical and quantum-mechanical tunneling times, expressed as $\tau_{\mathrm{QM}}=(\pi/2)\bar{\tau}$, where $\tau_{\mathrm{QM}}$ corresponds to half the oscillation period of the probability of finding the particle in either well. This relation is further confirmed for generic double-well systems through a WKB analysis. As a concrete application, the inversion dynamics of the ammonia molecule is analyzed, yielding an inversion frequency of approximately $24$ GHz, in close agreement with experimental observations. These results highlight the potential of stochastic quantization as a powerful and physically insightful framework for analyzing tunneling phenomena in quantum systems.
@@ I Introduction
Because time is not represented by an operator in quantum mechanics—but rather appears only as a parameter labeling the evolution of quantum states—the problem of defining, let alone measuring, the traversal time between two points for a quantum particle remains an elusive and, to some extent, controversial issue [HaugeStovnengRMP, LandauerMartinRMP1994, Hilgevoord2005, WinfulPhysRep2006, MugaBookVol1]. The question becomes particularly subtle in the context of quantum tunneling, where a particle traverses a potential barrier higher than its own energy—a process forbidden in classical dynamics. This raises fundamental questions, such as how long it takes for the particle to tunnel through the barrier or, conversely, how much time it spends within the classically forbidden region.
@@
In order to address these and related questions, a variety of operational and theoretical proposals for ‘traversal time’ have been advanced. The dwell time defines the average time spent by the particle in a specified spatial region, irrespective of whether it is ultimately transmitted or reflected [HaugeStovnengRMP]. The phase delay time extracts a time scale from the energy derivative of the scattering phase (or of the $S$-matrix), and is often applied to the transmitted component in one-dimensional tunneling [Wigner1955, Smith1960, DeMouraAlbuquerque_SSC_1990]. The Büttiker–Landauer (BL) time probes traversal by monitoring the system’s sensitivity to a slow modulation of the barrier, yielding an operational time linked to adiabatic response [BuettikerLandauerPRL, BuettikerPRB]. Larmor-clock times introduce a weak, localized magnetic field in the barrier and infer the time from spin precession [Buettiker1990, LandauerMartin1992]. Bohmian (trajectory-based) analyses likewise define traversal and residence times from guidance trajectories, offering a complementary, causal picture that can be directly compared with operational proposals [LeavensBohmian, OriolsMompart]. More recently, alternative formulations have sought to promote time to the status of a genuine quantum observable, such as the space–time-symmetric extensions of quantum mechanics that introduce a self-adjoint time operator and a dual, space-conditional Schrödinger equation within the same Hilbert-space framework [DiasParisioPRA2017, AraujoParisioPRA2024, FloresTOA2024, deLaraBeims2024_TravelingTime]. Despite this profusion of proposals, no single notion of traversal (or tunneling) time has gained universal acceptance: each highlights a particular operational or theoretical aspect of the process, yet each also entails limitations that preclude it from serving as a context-independent notion of “time” [HaugeStovnengRMP, LandauerMartinRMP1994, OlkhovskyRecamiPR2004, WinfulPhysRep2006, MugaBookVol1, TimeInQMVol2].
@@
Stochastic quantization (SQ) is a formalism introduced by Nelson in 1966 that describes quantum systems as conservative diffusion processes [Nelson1966]. In this framework, the dynamics of a quantum particle is governed by a stochastic differential equation whose deterministic drift term accounts for the interaction potential, while the diffusion coefficient (related to the strength of the underlying Brownian motion) is proportional to Planck’s constant. By imposing a stochastic analogue of Newton’s second law, Nelson demonstrated that stochastic quantization is mathematically equivalent to the Schrödinger equation. Subsequent developments  [YasueJFA1981, YasueJMP1981, GuerraMoratoPRD1983, PavonJMP1995, YangJMP2021] clarified the Lagrangian and Hamiltonian variational underpinnings of stochastic quantization and its unitary correspondence [Pavon2000] with conventional quantum mechanics. For both earlier and more recent reviews of stochastic quantization, see, e.g., Refs. [Guerra1981, ZambriniIJTP1985, Nelson1985, BlanchardCombeZheng1987, BeyerPaulUniverse2021, Kuipers2023].
@@
Because the SQ formalism is expressed in terms of real-space stochastic trajectories, it provides a natural framework for investigating dynamical quantities that are difficult to define or access within the standard quantum-mechanical formulation. One such quantity is precisely the traversal time between two spatial points. Indeed, by employing the concept of first-passage times from the theory of diffusion processes [Gardiner2009], one can evaluate not only the mean traversal time but also its full probability distribution—a quantity rarely addressed by other tunneling-time proposals discussed above. These aspects make stochastic quantization a powerful tool for studying tunneling phenomena, as demonstrated in previous works on tunneling through square barriers [ChenWang_PLA_1990, Hara2003, KudoNittaPLA2013] and on bound-state tunneling in a quartic double-well potential [KoeppeJMP2018_QHE_1D].
@@
In this paper, we employ stochastic quantization to investigate in detail the tunneling process in double-well potentials, where a particle passes through the barrier separating the two minima. We compute the mean tunneling time, denoted by $\bar{\tau}$, for each potential considered and analyze its dependence on relevant physical parameters. We show, in particular, that $\bar{\tau}$ has a clear physical interpretation, as it can be directly related to the corresponding—and spectroscopically measurable—“tunneling time” derived from standard quantum mechanics within the two-level approximation. We further show that the empirical distributions of tunneling times obtained from large ensembles of numerical trajectories exhibit exponential tails that are in excellent agreement with the theoretical predictions derived from first–passage time theory. Specifically, we first consider a square double-well potential, for which the mean tunneling time can be calculated analytically. Theoretical predictions are shown to be in excellent agreement with numerical simulations, thereby validating the analytical approach. We also analyze the high-barrier limit, establishing a direct relation between the tunneling time predicted by SQ and the corresponding tunneling time obtained from quantum mechanics (see below). We then demonstrate, through a WKB calculation, that this relation holds for generic double-well systems. Finally, we apply the SQ framework to the inversion problem of the ammonia molecule, in which the nitrogen atom oscillates between two equivalent configurations on opposite sides of the plane defined by the three hydrogen atoms—effectively tunneling through the potential barrier separating the two states.
@@
In quantum mechanics, tunneling in a double-well potential is typically treated within the two-state approximation. By constructing symmetric and antisymmetric superpositions of the ground and first excited states, one obtains states localized primarily in the left and right wells, respectively. If the system is initially prepared in one of these localized (nonstationary) states, the probability of finding the particle in either well oscillates periodically in time, with period $T$ inversely proportional to the energy splitting $\Delta E$. The oscillation half-period, $\tau_{\mathrm{QM}}={T}/{2}$, is commonly referred to as the quantum-mechanical tunneling time. We demonstrate—first analytically for the square double-well potential and subsequently through a WKB analysis for generic double-well systems—that, in the high-barrier limit, the following relation between the SQ mean tunneling time $\bar{\tau}$ and the quantum-mechanical tunneling time $\tau_{\mathrm{QM}}$ holds:
@@
This relation, previously observed numerically for the quartic double-well potential [BeyerPaulUniverse2021], is here derived analytically and shown to hold generally, provided the barrier is sufficiently high.
Having established these theoretical results, we next address the inversion problem of the ammonia molecule as a concrete tunneling example. To this end, we compute the tunneling times for a particle moving in a double Rosen–Morse potential, a model commonly employed for the ammonia inversion process [RosenMorse, YangHan_IJMS_2021_Ammonia]. Although this case cannot be solved analytically, the mean tunneling time can be obtained numerically by direct integration of the corresponding SQ expression. From the SQ mean tunneling time $\bar{\tau}$ and relation (1), we estimate the inversion frequency of the ammonia molecule. The resulting value—approximately $24$ GHz—is in excellent agreement with experimental observations, thus validating the SQ-based theoretical approach to tunneling-time calculations.
The remainder of this paper is organized as follows. Section II provides a brief review of stochastic quantization. In Sec. III, we discuss how the mean tunneling time and the full distribution of tunneling times can be obtained from the theory of first-passage times within the SQ framework. Section IV applies the SQ formalism to tunneling in a square double-well potential, where most calculations can be performed analytically, including the derivation of Eq. (1). Similar calculations are then extended in Sec. V, using a WKB approach to treat generic double-well systems. Section VI presents a concrete application of the SQ approach to the inversion problem of the ammonia molecule. The implications of our findings are discussed in Sec. VII, and Sec. VIII summarizes our main conclusions.
II Stochastic Quantization: Brief Review
Although introduced nearly six decades ago, Nelson’s formulation of stochastic quantization (also known as stochastic mechanics) is perhaps not widely known among physicists. For this reason, we include here a brief, self-contained exposition of its basic principles. Our treatment departs slightly from Nelson’s original presentation [Nelson1966] in that we demonstrate how one of Nelson’s dynamical conditions [see Eq. (6)] can be derived directly from the requirement of time reversibility, whereas in Nelson’s original formulation it was introduced as an ad hoc assumption.
In Nelson’s stochastic quantization approach, quantum states are described by the following pair of forward and backward stochastic differential equations:
@@ II.1 Hydrodynamical Equivalence to Quantum Mechanics
for some function $S(\mathbf{x},t)$. A related discussion is given in Ref. [delaPenaCetto2014], where it is shown that postulating that $\mathbf{v}$ is irrotational leads to Eq. (20). In contrast, here we have argued on physical grounds (time-reversal symmetry) that Eq. (20) must hold, which in turn implies that $\mathbf{v}$ is irrotational.
In Nelson’s original derivation, Eq. (22) was postulated as an assumption. This condition can also be justified within variational formulations of stochastic quantization [GuerraMoratoPRD1983, PavonJMP1995]. Now, using that $\mathbf{v}$ and $\mathbf{u}$ are both irrotational, Eq. (21) can be written as
@@
The remarkable discovery by Nelson [Nelson1966] was to note that by considering the function
@@
More precisely, with the identifications (27) and (28), Eqs. (23) and (25) correspond precisely to the two flow equations obtained by Madelung [Madelung1927] upon using the representation (26) in Schrödinger equation. In particular, with the identifications (27) and (26) the flow velocities $\mathbf{v}$ and $\mathbf{u}$ read
@@
In summary, Nelson’s stochastic quantization can be formulated in terms of two Newton-like dynamical laws that respect time-reversal symmetry, namely
@@
It is also possible to establish the unitary equivalence between Nelson’s stochastic quantization and quantum mechanics, which provides a rigorous mathematical foundation for the correspondence between the two formalisms (see Appendix). This shows that they are not merely analogous but are in fact related by a fundamental unitary transformation. This deep connection suggests that the stochastic and quantum perspectives are simply two different ways of looking at the same underlying physical reality.
@@ II.2 Tunneling Times: Mean and Distribution
Nelson’s quantization approach has the advantage that one can draw upon the theory of stochastic processes to investigate certain properties that are not easily accessible via standard quantum mechanics. One such quantity is the mean first passage time for a particle to exit a given interval. To be specific, let us consider the mean time that the particle takes to leave an interval $(a,b)$, starting from an initial point $x$. For our purposes here, let us assume that points $a$ and $b$ are reflecting and absorbing boundaries, respectively. Then, using standard formula from the theory of stochastic processes [Gardiner2009, PaulBaschnagel1999], the mean first passage time (MFPT) for the particle to cross the right boundary is given by
@@
The theory of first passage times enables us to study not only the MFPT, as shown above, but also the distribution of tunneling times, as discussed next. Let $f(x,t)$ be the probability that the particle will exit the interval $(a,b)$ at time $t$, after starting in $t=0$ from point $x$. Then one has [Gardiner2009, PaulBaschnagel1999]
@@
We will see in the remainder of the paper that both the mean tunneling time, given by Eq. (38), and the exponential tail behavior of the tunneling-time distribution, as predicted by Eq. (42), are in excellent agreement with results obtained from numerical simulations of Nelson’s stochastic process for specific potentials. (An exponential decay in the tunneling-time distribution was observed numerically for both the single-barrier case [Hara2003] and the quartic double-well potential [Koeppe2017], although apparently without recognition of the theoretical basis for this behavior, as established above.) As already mentioned, our main interest here concerns tunneling processes in double-well potentials that bear direct application, for instance, to the problem of inversion oscillation of the ammonia molecule. We begin by considering the double square-well potential, for which analytical calculations can be carried out.

### v2

# Tunneling in double-well potentials within Nelson’s stochastic mechanics: Application to ammonia inversion

@@ Abstract
Nelson’s stochastic mechanics formulates quantum dynamics as a real-time conservative diffusion process in which a particle undergoes Brownian-like motion with a fluctuation amplitude fixed by Planck’s constant. While being mathematically equivalent to the Schrödinger formulation, this approach provides an alternative dynamical framework that enables the study of time-resolved quantities that are not straightforwardly defined within the standard operator-based approach. In the present work, Nelson’s stochastic mechanics is employed to investigate tunneling-time statistics for bound states in double-well potentials. Using first-passage time theory within this framework, both the mean tunneling time, $\bar{\tau}$, and the full probability distribution, $p({\tau})$, are computed. The theoretical predictions are validated through extensive numerical simulations of stochastic trajectories for two representative potentials. For the square double-well potential, analytical expressions for $\bar{\tau}$ are derived and are shown to be in excellent agreement with simulations. In the high-barrier limit, the results reveal a direct relation between the stochastic-mechanical and quantum-mechanical tunneling times, expressed as $\tau_{\mathrm{QM}}=(\pi/2)\bar{\tau}$, where $\tau_{\mathrm{QM}}$ corresponds to half the oscillation period of the probability of finding the particle in either well. This relation is further confirmed for generic double-well systems through a WKB analysis. As a concrete application, the inversion dynamics of the ammonia molecule is analyzed, yielding an inversion frequency of approximately $24$ GHz, in close agreement with experimental observations. These results highlight the potential of stochastic mechanics as a conceptually coherent and quantitatively consistent framework for analyzing tunneling phenomena in quantum systems.
@@ I Introduction
Because time is not represented by an operator in quantum mechanics—but rather appears only as a parameter labeling the evolution of quantum states—the problem of defining, let alone measuring, the traversal time between two points for a quantum particle remains an elusive and, to some extent, controversial issue [23, 32, 24, 63, 41]. The question becomes particularly subtle in the context of quantum tunneling, where a particle traverses a potential barrier higher than its own energy—a process forbidden in classical dynamics. This raises fundamental questions, such as how long it takes for the particle to tunnel through the barrier or, conversely, how much time it spends within the classically forbidden region.
@@
In order to address these and related questions, a variety of operational and theoretical proposals for ‘traversal time’ have been advanced. The dwell time defines the average time spent by the particle in a specified spatial region, irrespective of whether it is ultimately transmitted or reflected [23]. The phase delay time extracts a time scale from the energy derivative of the scattering phase (or of the $S$-matrix), and is often applied to the transmitted component in one-dimensional tunneling [62, 58, 14]. The Büttiker–Landauer (BL) time probes traversal by monitoring the system’s sensitivity to a slow modulation of the barrier, yielding an operational time linked to adiabatic response [6, 7]. Larmor-clock times introduce a weak, localized magnetic field in the barrier and infer the time from spin precession [8, 31]. Bohmian (trajectory-based) analyses likewise define traversal and residence times from guidance trajectories, offering a complementary, causal picture that can be directly compared with operational proposals [33, 49]. More recently, alternative formulations have sought to promote time to the status of a genuine quantum observable, such as the space–time-symmetric extensions of quantum mechanics that introduce a self-adjoint time operator and a dual, space-conditional Schrödinger equation within the same Hilbert-space framework [17, 1, 18, 13]. Despite this profusion of proposals, no single notion of traversal (or tunneling) time has gained universal acceptance: each highlights a particular operational or theoretical aspect of the process, yet each also entails limitations that preclude it from serving as a context-independent notion of “time” [23, 32, 48, 63, 41, 42].
@@
Stochastic mechanics (SM) is a formalism introduced by Nelson in 1966 that describes quantum systems as conservative diffusion processes [43]. In Nelson’s framework, the dynamics of a quantum particle is governed by a stochastic differential equation whose deterministic drift term accounts for the interaction potential, while the diffusion coefficient (related to the strength of the underlying Brownian motion) is proportional to Planck’s constant. By imposing a stochastic analogue of Newton’s second law, Nelson demonstrated that stochastic mechanics is mathematically equivalent to the Schrödinger equation. Subsequent developments  [67, 66, 20, 53, 65] clarified the Lagrangian and Hamiltonian variational underpinnings of stochastic mechanics and its unitary correspondence [52] with conventional quantum mechanics. For both earlier and more recent reviews of stochastic mechanics, see, e.g., Refs. [21, 68, 45, 5, 3, 29]. Nelson’s formalism may also be viewed, in hindsight, as a dynamical counterpart of Schrödinger bridge problem introduced in 1931 [56]. In that framework, one selects—via a relative-entropy minimization—the most probable diffusion process interpolating between prescribed endpoint distributions. This problem is now understood as a precursor to modern mass-transport theory [34, 40]; see also Ref. [10] for an English translation of Schrödinger’s original paper and a detailed discussion of its connections to large-deviation theory and stochastic optimal control.
It should be noted that the formal equivalence between SM and the Schrödinger equation has been critically examined in the literature. In particular, Wallstrom [60, 61] observed that formulations expressed solely in terms of the hydrodynamic variables—namely, the probability density and the phase (see below)—may admit solutions that do not automatically satisfy the single-valuedness condition of the Schrödinger wavefunction, thereby indicating that an additional quantization condition on the phase may be required. Subsequent analyses have argued that appropriate regularity, topological, or variational considerations remove the spurious solutions; see, e.g., Ref. [29]. In the present work we restrict our attention to states corresponding to standard Schrödinger solutions, and thus operate entirely within the conventional quantum sector of the theory.
A related concern, already emphasized by Nelson [45], pertains to the apparent mismatch between multi-time correlations of the underlying diffusion process and the operator-ordered correlation functions of textbook quantum theory. This issue has been analyzed in detail by Blanchard et al. [4], who showed that a proper treatment of measurement-induced state updating within the stochastic framework restores agreement with standard quantum predictions; more recent discussions have further clarified this point within both operational and modified dynamical formulations of SM [16, 29]. Nelson also emphasized that, for entangled systems, the stochastic drift in configuration space exhibits explicit nonlocal dependence [46]. From a contemporary perspective, however, this feature is not generally regarded as a defect but rather as a structural property shared by realist trajectory-based reconstructions of quantum mechanics that successfully reproduce entanglement phenomena. In this sense, the nonlocal character of SM may be viewed as reflecting its compatibility with the implications of Bell-type results rather than signaling an internal inconsistency of the theory; see also recent discussions in Refs. [16, 29].
@@
Because the SM formalism is expressed in terms of real-space stochastic trajectories, it provides a natural framework for investigating dynamical quantities that are difficult to define or access within the standard quantum-mechanical formulation. One such quantity is precisely the traversal time between two spatial points. Indeed, by employing the concept of first-passage times from the theory of diffusion processes [19], one can evaluate not only the mean traversal time but also its full probability distribution—a quantity rarely addressed by other tunneling-time proposals discussed above. These aspects make stochastic mechanics a powerful tool for studying tunneling phenomena, as demonstrated in previous works on tunneling through square barriers [9, 22, 28] and on bound-state tunneling in a quartic double-well potential [26].
@@
In this paper, we employ Nelson’s stochastic mechanics to investigate in detail the tunneling process in double-well potentials, where a particle passes through the barrier separating the two minima. We compute the mean tunneling time, denoted by $\bar{\tau}$, for each potential considered and analyze its dependence on relevant physical parameters. We show, in particular, that $\bar{\tau}$ has a clear physical interpretation, as it can be directly related to the corresponding—and spectroscopically measurable—“tunneling time” derived from standard quantum mechanics within the two-level approximation. We further show that the empirical distributions of tunneling times obtained from large ensembles of numerical trajectories exhibit exponential tails that are in excellent agreement with the theoretical predictions derived from first–passage time theory. Specifically, we first consider a square double-well potential, for which the mean tunneling time can be calculated analytically. Theoretical predictions are shown to be in excellent agreement with numerical simulations, thereby validating the analytical approach. We also analyze the high-barrier limit, establishing a direct relation between the tunneling time predicted by SM and the corresponding tunneling time obtained from quantum mechanics (see below). We then demonstrate, through a WKB calculation, that this relation holds for generic double-well systems. Finally, we apply the SM framework to the inversion problem of the ammonia molecule, in which the nitrogen atom oscillates between two equivalent configurations on opposite sides of the plane defined by the three hydrogen atoms—effectively tunneling through the potential barrier separating the two states.
@@
In quantum mechanics, tunneling in a double-well potential is typically treated within the two-state approximation. By constructing symmetric and antisymmetric superpositions of the ground and first excited states, one obtains states localized primarily in the left and right wells, respectively. If the system is initially prepared in one of these localized (nonstationary) states, the probability of finding the particle in either well oscillates periodically in time, with period $T$ inversely proportional to the energy splitting $\Delta E$. The oscillation half-period, $\tau_{\mathrm{QM}}={T}/{2}$, is commonly referred to as the quantum-mechanical tunneling time. We demonstrate—first analytically for the square double-well potential and subsequently through a WKB analysis for generic double-well systems—that, in the high-barrier limit, the following relation between the SM mean tunneling time $\bar{\tau}$ and the quantum-mechanical tunneling time $\tau_{\mathrm{QM}}$ holds:
@@
This relation, previously observed numerically for the quartic double-well potential [3], is here derived analytically and shown to hold generally, provided the barrier is sufficiently high.
Having established these theoretical results, we next address the inversion problem of the ammonia molecule as a concrete tunneling example. To this end, we compute the tunneling times for a particle moving in a double Rosen–Morse potential, a model commonly employed for the ammonia inversion process [55, 64]. Although this case cannot be solved analytically, the mean tunneling time can be obtained numerically by direct integration of the corresponding SM expression. From the SM mean tunneling time $\bar{\tau}$ and relation (1), we estimate the inversion frequency of the ammonia molecule. The resulting value—approximately $24$ GHz—is in excellent agreement with experimental observations, thus validating the SM-based theoretical approach to tunneling-time calculations.
The remainder of this paper is organized as follows. Section II provides a brief review of stochastic mechanics. In Sec. III, we discuss how the mean tunneling time and the full distribution of tunneling times can be obtained from the theory of first-passage times within the SM framework. Section IV applies the SM formalism to tunneling in a square double-well potential, where most calculations can be performed analytically, including the derivation of Eq. (1). Similar calculations are then extended in Sec. V, using a WKB approach to treat generic double-well systems. Section VI presents a concrete application of the SM approach to the inversion problem of the ammonia molecule. The implications of our findings are discussed in Sec. VII, and Sec. VIII summarizes our main conclusions.
II Stochastic Mechanics: Brief Review
Although introduced nearly six decades ago, Nelson’s formulation of stochastic mechanics (also known as stochastic quantization) is perhaps not widely known among physicists. For this reason, we include here a brief, self-contained exposition of its basic principles. Our treatment departs slightly from Nelson’s original presentation [43] in that we demonstrate how one of Nelson’s dynamical conditions [see Eq. (6)] can be derived directly from the requirement of time reversibility, whereas in Nelson’s original formulation it was introduced as an ad hoc assumption.
In Nelson’s stochastic mechanics approach, quantum states are described by the following pair of forward and backward stochastic differential equations:
@@ II.1 Hydrodynamical Equivalence to Quantum Mechanics
for some function $S(\mathbf{x},t)$. A related discussion is given in Ref. [12], where it is shown that postulating that $\mathbf{v}$ is irrotational leads to Eq. (20). In contrast, here we have argued on physical grounds (time-reversal symmetry) that Eq. (20) must hold, which in turn implies that $\mathbf{v}$ is irrotational.
In Nelson’s original derivation, Eq. (22) was postulated as an assumption. This condition can also be justified within variational formulations of stochastic mechanics [20, 53]. Now, using that $\mathbf{v}$ and $\mathbf{u}$ are both irrotational, Eq. (21) can be written as
@@
The remarkable discovery by Nelson [43] was to note that by considering the function
@@
More precisely, with the identifications (27) and (28), Eqs. (23) and (25) correspond precisely to the two flow equations obtained by Madelung [36] upon using the representation (26) in Schrödinger equation. In particular, with the identifications (27) and (26) the flow velocities $\mathbf{v}$ and $\mathbf{u}$ read
@@
In summary, Nelson’s stochastic mechanics can be formulated in terms of two Newton-like dynamical laws that respect time-reversal symmetry, namely
@@
It is also possible to establish the unitary equivalence between Nelson’s stochastic mechanics and standard quantum mechanics, which provides a rigorous mathematical foundation for the correspondence between the two formalisms (see Appendix). This shows that they are not merely analogous but are in fact related by a fundamental unitary transformation. This deep connection suggests that the stochastic and quantum perspectives are simply two different ways of looking at the same underlying physical reality.
@@ II.2 Tunneling Times: Mean and Distribution
Nelson’s quantization approach has the advantage that one can draw upon the theory of stochastic processes to investigate certain properties that are not easily accessible via standard quantum mechanics. One such quantity is the mean first passage time for a particle to exit a given interval. To be specific, let us consider the mean time that the particle takes to leave an interval $(a,b)$, starting from an initial point $x$. For our purposes here, let us assume that points $a$ and $b$ are reflecting and absorbing boundaries, respectively. Then, using standard formula from the theory of stochastic processes [19, 51], the mean first passage time (MFPT) for the particle to cross the right boundary is given by
@@
The theory of first passage times enables us to study not only the MFPT, as shown above, but also the distribution of tunneling times, as discussed next. Let $f(x,t)$ be the probability that the particle will exit the interval $(a,b)$ at time $t$, after starting in $t=0$ from point $x$. Then one has [19, 51]
@@
We will see in the remainder of the paper that both the mean tunneling time, given by Eq. (38), and the exponential tail behavior of the tunneling-time distribution, as predicted by Eq. (42), are in excellent agreement with results obtained from numerical simulations of Nelson’s stochastic process for specific potentials. (An exponential decay in the tunneling-time distribution was observed numerically for both the single-barrier case [22] and the quartic double-well potential [27], although apparently without recognition of the theoretical basis for this behavior, as established above.) As already mentioned, our main interest here concerns tunneling processes in double-well potentials that bear direct application, for instance, to the problem of inversion oscillation of the ammonia molecule. We begin by considering the double square-well potential, for which analytical calculations can be carried out.